OpenStudy (anonymous):

There are 8 brooms and 6 mops in a janitor's closet. What is the ratio of the number of mops to the total number of brooms and mops? A: 4/3 B:3/7 C:7/3 D:3/4

OpenStudy (anonymous):

helppp!!!.

OpenStudy (anonymous):

Mops/brooms= 6/8 =3/4 we will just simplify to its lowest term :)

OpenStudy (anonymous):

Mops =6, Total= 8+6=? Ratio =?

OpenStudy (anonymous):

@saloniiigupta95 i don't think so

OpenStudy (anonymous):

Think what??? *ratio of the number of 'mops' to the 'total number of brooms and mops'?* makes sense to me MOPS/ TOTAL OF MOPS AND BROOMS ...

OpenStudy (anonymous):

Ohh yeahh my mistake sorry :( Mops/brooms+mops=6/6+8 =6/14=3/7 So, its option B

OpenStudy (anonymous):

Yups.... no worries :-)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

actualy your wrong, its 3/4